# SVD-Free Low-Rank Adaptive Gradient Optimization for Large Language Models

### v1

@@ 5.2 DCT as linear approximation of the gradient eigenbasis
6 Experiments
We now present our numerical results. Our main goal is to show that DCT projection can recover the performance of SVD-based projection and have lower memory footprint and running time.
For pre-training, we use a Llama model (Touvron et al., 2023) with 800 million parameters and train it on the C4 dataset using 20/100 tokens/parameter, equivalent to 16/80B tokens. We replace the SVD projection in Frugal and Fira with DCT matrix to directly compare the optimizer behavior when the only difference is the projection. We use a learning rate tuned for the full-rank AdamW baseline to avoid the expensive learning rate tuning process for each individual projection. As a result, in certain cases, the DCT projection would require further learning rate tuning. In all settings, our main comparison is against the SVD-like projection and we will also add the full-rank AdamW results for reference. For fine-finetuning, we use Llama-2 7B and report the accuracy on GSM-8k dataset (Cobbe et al., 2021) using Frugal and Fira with SVD and DCT projections, as well as on LDAdamW and our DCT-AdamW implementation.
6.1 Pre-training (PT)
We train the Llama-800M model from scratch using 16B and 80B tokens, equivalent to 20 and 100 tokens per parameter, from C4 dataset (Raffel et al., 2020). The hidden dimension of the model is $2048$ and we use rank $r=256$ for all runs, as well as learning rate $\eta=7.5e-5$, which was chosen for the full-rank AdamW, proportionally to the inverse square root of the number of parameters and kept fixed for all runs.
@@ 6.1 Pre-training (PT)
PT with FRUGAL. We integrate the DCT matrix into the FRUGAL optimizer and compare against the SVD, RandPerm and Random projections. RandPerm uses a random permutation as a projection matrix, while Random generates a random, semi-orthogonal matrix. In Figure 1(a) we show the training loss for all these runs. In the zoomed-in plot for the last 5000 training steps we see that SVD projection recovers the performance full-rank AdamW, while the DCT projection is a good approximation of the SVD, which supports our claim. We present our numerical results in Table 3. We would like to emphasize the running time reduced by 1h 48m ($\approx 22.6\%$) compared to the SVD projection, as well as the memory usage reduced by about 2.2GB ($\approx 3.5\%$), while the increase in perplexity is less than one point. In comparison to RandPerm and Random projections, the runtime is on par, while train and validation perplexities are lower by approximately one point in the favor of DCT, illustrating again the benefit of our approach.
@@
PT with FIRA. We integrate DCT into FIRA optimizer and compare against SVD. In Figure 1(b) we show the training loss, where our focus is on the comparison between SVD and DCT projections. We observe that DCT consistently yields lower training loss compared to the SVD projection, which is also visible in the numerical results in Table 3, translated to lower perplexities in the favor of DCT. Moreover, the memory usage is smaller by 2GB ($\approx 3\%$) and the running time is lower by 1h 54m ($\approx 23.8\%$).
@@
[…]
AdamW LDAdamW DCT-AdamW
Train PPL 12.88 15.10 14.95
Val. PPL 11.73 13.91 13.69
Mem. (GiB) 73.72 72.10 57.82
Time 1d 13h 22m 2d 1h 24m 1d 15h 17m
6.2 Fine-tuning (FT)
We finetune Llama-2 7B on GSM-8k dataset using SVD and DCT projections for FRUGAL and FIRA optimizers, as well as LD-AdamW and DCT-AdamW. We show our results for ranks $r\in\{32,512\}$ in Table 5.
FT with FRUGAL. Both projections yield similar training loss for both rank values, which is an indication that DCT is indeed a good approximation for SVD. Despite having different training loss, the SVD projection achieves the same accuracy for both ranks. It is surprising that DCT yields better accuracy for lower rank compared to higher rank. In terms of memory usage, as expected for this large model, the DCT projection saves 8GB of memory ($\approx 23.4\%$) for $r=32$, while for $r=512$ the reduction in memory is only 6.3GB ($\approx 18.2\%$). The running time is reduced by roughly 35m ($\approx 75\%$) for both ranks.

### v2

@@ 4.2 DCT as linear approximation of the gradient eigenbasis
5 Related Work
Our approach aims to improve existing optimizers from prior work in the literature, which we group into two categories: optimizers that speed up convergence at the expense of increased FLOPs and optimizers focused on reducing the running time and memory usage of optimizers by compressing the gradient via low-rank matrix factorization.
Fast Convergence Optimizers. Muon optimizer (Jordan et al., 2024) has recently stood up in the literature for its fast convergence rate due to the orthogonalized momentum update. The purpose of the orthogonalization is to push the singular values of the momentum matrix towards $1$. It speeds up convergence by increasing the importance (singular values) of directions in the momentum matrix, which otherwise would have a low impact over the optimization.
The most straightforward approach to computing an orthogonal update is to use the $UV^{\top}$ from the SVD decomposition of the momentum matrix. Since SVD is expensive, it is replaced by an iterative procedure called Newton-Schulz Newton-Schulz, which involves computing the odd powers of the momentum matrix up to $5^{th}$ order and multiplying by some carefully chosen constants. While the Newton-Schulz procedure delivers an accurate approximation of $UV^{\top}$, the odd powers in the polynomial involve full-size matrix multiplications. This is in particular difficult for large scale settings, where the full matrices have to be materialized on a GPU before running Newton-Schulz, which increases communication and memory usage. Dion optimizer (Ahn et al., 2025) aims to reduce the communication overhead by using low-rank, orthogonal updates computed via Power-Iteration, a procedure that requires QR decomposition.
Instead, we propose to reduce the overhead of Newton-Schulz in Muon and QR-decomposition in Dion by factorizing the momentum to a low-rank matrix using our DCT-based dynamic column selection approach and orthogonalizing the low-rank momentum using Newton-Schulz, then project back to the original space to update the model parameters.
Low-Rank Compression of Optimizer States. Most approaches use individual projection matrices tailored to the gradient at each layer by invoking techniques based on quantization or matrix factorization, such as SVD, QR or PCA to reduce the memory usage of the optimizer states stored in GPU memory. The most common approach is to factorize the gradient to low-rank matrices. One pioneering work in the context of low-rank adaptive optimization is the recent GaLore optimizer (Zhao et al., 2024), which performs SVD once at a few steps to project the gradient to a lower-dimensional space. GaLore was followed by several other optimizers that improve certain aspects of it. LDAdam (Robert et al., 2025) aims to improve the computational runtime of GaLore by compressing the first order momentum in AdamW, replacing SVD with a block power-iteration (Bentbib & Kanber, 2015) and performing a smooth subspace transition by rotating the first and second momentum accordingly to incorporate gradients from the same subspace at each step.
By default, GaLore discards the projection error, which LDAdam stores and incorporates into the gradient at the next step to improve convergence. In contrast, FRUGAL (Zmushko et al., 2024) makes the distinction between the low-rank gradient (called state-full) and the projection error (called state-free). The state-full gradient is used in AdamW, while the state-free gradient is fed to an optimizer without a state, such as SignSGD. The main idea is to leverage the remaining gradient information in the projection error since it is available at each step instead of discarding or storing it. A concurrent work to FRUGAL is FIRA (Chen et al., 2024), which preserves the low-rank constraint for memory efficiency while achieving full-rank performance by properly scaling the projection error.
Another approach worth mentioning is Online Subspace Descent Liang et al. (2024), which replaces SVD projection with Online PCA and involves computing the projection matrix as a solution of an optimization objective focused on: (a) minimizing the projection error and (b) forcing the projection matrix to be orthogonal. The authors indicate that performing one step with Adam to solve this additional optimization problem is enough to obtain a qualitatively good projection matrix $P$. In contrast, our method does not introduce such overheads during training since the DCT matrix is already computed at the beginning of training.
Despite all aforementioned approaches being similar in using SVD/QR-based projections, they differ in a few aspects. The most important one in our view is how they handle the projection error and how often they update the low-dimensional subspace, which we clarify in Table 3.
Our approach uses DCT projection coupled with a dynamic column selection to determine a projection matrix tailored to the gradient/momentum for a particular layer and can be integrated into any optimizer, regardless of the way it handles the projection error.
Low-rank Projection Type Frequency* Error
GaLore (Zhao et al., 2024) SVD 200 discard
FIRA (Chen et al., 2024) SVD 200 norm-based scaling
LDAdam (Robert et al., 2025) Block Power Iteration 1 error feedback
Dion (Ahn et al., 2025) Power Iteration 1 save to momentum
Trion (this work) DCT any same as Dion
DCT-AdamW (this work) DCT any error feedback
6 Conclusion and Limitations
We introduced Trion and DCT-AdamW optimizers that use our DCT-based dynamic column selection to replace two techniques used to perform low-rank decomposition, the inaccurate Power-Iteration in Dion and the expensive SVD in FRUGAL/FIRA and showed that our work improves running time, memory usage and recovers the accuracy of the original methods and thus serves as a cheaper alternative to these expensive and inaccurate methods used to perform low-rank decomposition in adaptive gradient methods for both pretraining and finetuning.
Our experiments are limited to models with at most 1.3B parameters for pretraining and additional work is required to test our technique for larger models and beyond the Chinchilla-optimal token counts, which would require significantly more computational resources.
Appendix A Efficient Computation of the DCT Matrix on GPU
In this section we present how we can efficiently compute the DCT-III matrix on a GPU using a vectorized implementation that is fast for large values of $n$ on the GPU. To create the DCT-III matrix $Q\in\mathbb{R}^{n\times n}$, we first create one vector $L=[0,\cdots,n-1]^{\top}$, which is used to create the matrix $\mathcal{I}\in\mathbb{N}^{n\times n}$ by replicating $L$ on the columns of $\mathcal{I}$:
We restate that we need to divide the first row by $\sqrt{2}$ to make $Q$ orthogonal. The computational efficiency comes from the element-wise product $\mathcal{I}\odot(2\mathcal{I}^{\top}+1)$ that actually computes the integer entries $i(2j+1)$. The DCT-II matrix can be obtained by transposing the DCT-III.
Appendix B Dynamic Column Selection Approach
In this section we provide further details about our dynamic column selection approach. When we compute the similarities $S$, we look at the columns of $S$ where on $i^{th}$ column we have the scalar products between all rows of $G$ and the $i^{th}$ column of $Q$. By choosing the columns with largest $\ell_{1}$- or $\ell_{2}$-norms we make sure we select the columns with largest overall alignment with all rows in $G$.
[…]
Appendix C Motivation of Discrete Cosine Transform Matrix
In general, any orthogonal matrix would yield similar quantitative results when used with our dynamic column selection technique. In this section we are interested in exploring different types of orthogonal matrices and our goal is to find a candidate matrix whose structure allows us to reduce the computational complexity for the operation $S=GQ$, especially for large layers.
The first candidate is a random orthogonal matrix $Q$, which can be obtained by generating a random Gaussian matrix and then orthogonalizing it using QR-decomposition (e.g. keeping only the Q-component from the decomposition), which should be done only once at the beginning of training. While this would work, it has the drawback of always having $O(n^{3})$ complexity when computing the similarities $S$ because the matrix does not have any structure that would allow us to use an algorithm with lower complexity.
The second candidate is the Hadamard matrix, a particular Fourier matrix containing only values $\pm 1$, which is used in the model compression literature and is known for its fast multiplication routines tailored to GPUs. However, it has the drawback of being ill defined for certain values of model’s dimension $d_{\rm model}$ and the current existing procedures provide a matrix that is not orthogonal in these cases, making it unusable for our context.
Our preferred candidate is the DCT matrix, which is also a Fourier matrix and has a particular structure that gives it the potential of computing $S=GQ$ ($S$ is called the discrete cosine transform of $G$), making it faster than the $O(n^{3})$ complexity of matmul, as we present below.
When $Q$ is the DCT-II matrix, we can reduce the complexity of $S=GQ$ from $O(n^{3})$ to $O(n^{2}log(n))$ by using the Makhoul’s $N$-point algorithm Makhoul (1980) to compute DCT faster. We summarize the algorithm in Appendix D and also provide a benchmark for different layer sizes and data types encountered in practice. In short, this is a FFT-based algorithm that benefits from the pre- and post-processing steps of the input matrix $G$ (such as permutations and multiplications with complex values) to reduce the computational complexity. Interestingly, the link between Makhoul’s algorithm to compute the cosine transform and the same transformation computed via a basic matmul is that the matmul version embeds all operations of the Makhoul’s algorithm in the DCT-II matrix itself, but at a higher computational cost.
Given the current data types implemented in PyTorch, Makhoul’s algorithm can be run only on float32 inputs, which makes it infeasible for practical bfloat16 inputs unless the input size is a power of 2. The limitation comes from the lack of complex-bfloat16 type support in PyTorch to represent the real and imaginary parts of a complex number as bfloat16. See Appendix D for more details.
Appendix D Makhoul’s Algorithm Explained
Makhoul’s $N$-point algorithm to compute a fast DCT can be summarized as follows:
  1. 1.
    apply a permutation to the input signal $X$ to obtain $X_{PERM}$: vector $[a,b,c,d,e,f]$ becomes $[a,c,e,f,d,b]$, where odd indices are in increasing order, even indices are in decreasing order and they are interleaved (can be cached for the same input size);
  2. 2.
    compute FFT of the permuted signal $X_{PERM}$ and obtain $X_{FFT}$;
  3. 3.
    compute Fourier coefficients $W_{k}=exp(-2i\pi k/N)$ (can be cached for the same input size);
  4. 4.
    obtain DCT of $X$ as $X_{DCT}=Real(X_{FFT}*W)$ (multiplication by $W$ accounts for permuting the input $X$).
Makhoul’s algorithm performs a one-dimensional, type-II DCT for each row of matrix $G$. In essence, the Makhoul’s algorithm $S=\textsc{Makhoul}(G)$ is equivalent to the matmul version $S=GQ$, where the matrix $Q$ embeds all the operations performed by the Makhoul’s procedure.
We benchmark Makhoul’s algorithm against the matmul version on GPU for different layer sizes and show it is faster than matmul for large layers. For small layers (up to embedding size $2048$, the benefits of Makhoul’s algorithm do not have practical benefits compared to matmul - they have similar runtimes). Concretely, we run 10 iterations of warmup before starting to measure the time for 100 runs on the GPU for both matmul version and Makhoul’s version, where $G$ is initialized with random gaussian data in float32 and stays fixed during the benchmark.
In Table 4 we show our benchmarking results where gradient matrix $G$ is in float32 format, compare against the DCT transform obtained via matmul by simply employing $S=GQ$ and compare against the Makhoul’s N-point algorithm. The column Ratio shows how much faster the Makhoul’s implementation is compared to the standard matmul. We can get up to $50\times$ speedup for matrices with more columns than rows.
Input size Source Model Matmul time (s) Makhoul time (s) Ratio (@ / FFT)
[…]
Table 4: Comparison of Matmul vs. Makhoul runtimes across different input sizes and models for float32. Ratio measures how much faster the Makhoul’s algorithm is compared to the standard matmul. Matmul is faster for ratio <1<1 and Makhoul is faster when ratio >1>1.
While we can get a significant speedup for the case $R<C$ for float32, the real practical settings use bfloat16 instead of float32 and for this reason we also run our benchmark for bfloat16 as follows: we generate a random matrix $G$ in float32 which will be used for Makhoul’s algorithm and the same matrix $G$ will be converted to bfloat16. Moreover, the DCT matrix $Q$ will also be stored in bfloat16 such that both operands in $S=GQ$ are in bfloat16. On the other hand, since Makhoul’s algorithm uses FFT from PyTorch, we are currently restricted to using float32 inputs because at the moment of developing this work PyTorch does not have the complex-bfloat16 type, where both real and imaginary parts are stored in bfloat16 (half precision is supported only for inputs with sizes power of two). As a result, we run Makhoul’s algorithm only in float32 and compare with matmul in bfloat16.
In Table 5 we show our benchmarking results where we compare matmul run with bfloat16 inputs against Makhoul’s algorithm run on float32 input. For $R\geq C$, matmul-bfloat16 is consistently faster than Makhoul’s algorithm. This is expected because bfloat16 type has higher throughput on GPUs than float32. However, we see that Makhoul’s algorithm is $3.5\times$ faster than matmul-bfloat16 for $R<C$, which is a much lower speedup compared to the results in Table 4 for float32.
Input size Source Model Matmul time (s) Makhoul time (s) Ratio (@ / FFT)
(4096, 4096) Llama-2-7B 0.00018977 0.00034137 0.56x
(25600, 5120) Qwen3-32B 0.00184539 0.00268176 0.69x
(5120, 25600) Qwen3-32B 0.00968907 0.00273731 3.54x
Table 5: Comparison of Matmul (bfloat16) vs. Makhoul (float32) timings across different input sizes and models with different data types.
The FFT-based algorithm shows reduction in the running time in our benchmark, as the theory predicts (e.g. $O(n^{2}log(n))$ compared to $O(n^{3})$) for float32, while the gains are limited when we compare against matmul-bfloat16. In order to benefit from the faster computation of Makhoul’s algorithm, we would need to convert bfloat16 matrices to float32, run the faster procedure, then convert back to bfloat16, which is not feasible because of the additional memory and computational overhead. However, when running training in mixed precision, the gradients are computed in bfloat16 and accumulated in a float32 buffer for precision reasons. Fortunately, we have access to this float32 buffer in the optimizer’s step function in PyTorch.
Appendix E Pseudocode of DCT-AdamW Optimizer
In this section we present the pseudocode of the DCT-AdamW optimizer, which uses our DCT-based dynamic column selection approach to create a dynamic, low-rank projection matrix use to factorize the gradient to a low-rank matrix.
Following LDAdamW (Robert et al., 2025), DCT-AdamW incorporates low-rank gradients from different subspaces and this requires rotating the momentum buffers using a rotation matrix $R$, where the second momentum buffer $v_{t}$ is updated using a simpler rule compared to LDAdamW. DCT-AdamW also supports error feedback and it saves only the error induced by the low-rank compression.
Algorithm 2 DCT-AdamW (right projection)
1:Input: $\beta_{1},\beta_{2},\epsilon,T,T_{u},r$
[…]
6:for $t=\{1,2,\dots,T\}$ do
7:  $G_{t}\leftarrow\nabla_{\theta}f(\theta_{t})+\Xi_{t}$
8:  $R\leftarrow\textsc{UpdateSubspace}(G_{t})$
9:  $g_{t}\leftarrow G_{t}\cdot Q_{crt}$ $\diamond$ projected gradient
10:  $\Xi_{t}\leftarrow G_{t}-g_{t}\cdot Q_{crt}^{\top}$ $\diamond$ update EF
11:  $m_{t}\leftarrow\beta_{1}\cdot m_{t-1}\cdot R+(1-\beta_{1})g_{t}$
12:  $v_{t}\leftarrow\beta_{2}|v_{t-1}\cdot R|+(1-\beta_{2})g_{t}^{2}$
14:end for
Algorithm 3 Update Subspace procedure
1:procedure UpdateSubspace($G$)
2:  $R\leftarrow I_{r\times r}$
3:  if $t>1$ then
4:   $\mathcal{I}_{prev}\leftarrow\mathcal{I}_{crt}$
5:  end if
6:  if $(t=1)\lor(t\mod T_{u}=0)$ then
7:   $S=\textsc{Makhoul}(G)$ $\triangleright$ or $S=GQ$
8:   $\mathcal{I}_{crt}\leftarrow\textsc{RankCols}(GQ,r)$
9:   $R\leftarrow Q_{prev}^{\top}\cdot Q_{crt}$
10:  end if
11:  return $R$
12:end procedure
The sets $\mathcal{I}_{prev/crt}$ hold the indices of the $r$ columns for the previous/current projections and $Q_{prev/crt}$ contain the columns from $Q$ specified by the these two sets. $T_{u}$ represents the subspace update interval (set to 200 for GaLore and to $1$ for LDAdam). The procedure RankCols ranks the columns of $S$, which is computed as $S=\textsc{Makhoul}(G)$ (see Appendix D) or $S=G_{t}\cdot Q$ as explained in Section 2.1. In order to make sure the momentum buffers integrate gradients from the same subspaces, we need to rotate $m_{t}$ and $v_{t}$ using a rotation matrix $R$ that first projects the momentum to the full-dimensional space using the previous projector and then projects it back to the current lower-dimensional subspace. We can perform this rotation directly in the $r$-dimensional space by multiplying the two projection matrices, resulting in $R\in\mathbb{R}^{r\times r}$. Rotating $v_{t}$ might introduce negative values and we apply the absolute value function to force the non-negativity of $v_{t}$. The low-rank gradient $g_{t}$ is computed by multiplying the full-rank gradient $G_{t}$ with $Q_{crt}$ and the full-rank update is obtained multiplying $u_{t}$ by $Q_{crt}^{\top}$.
Appendix F Projection Errors of Trion and Dion
We train two Llama models with 30M parameters ($d_{\rm model}=640$) with rank $r=128$ using Dion and Trion optimizers and save the gradients. For each set of gradients we simulate the operations of the other optimizer and record the metrics $\Delta_{t}^{\rm dion}$ and $\Delta_{t}^{\rm trion}$. Our experiments showed no significant differences in the patterns for the two simulations and we plot the projection errors for the simulation that used gradients from the model trained with Dion optimizer. In Figure 1 we show the projection errors for the linear layers in the first transformer block, where Trion yields lower projection error than Dion and the trend is the same across all transformer blocks in the model. These plots explain why Trion has lower training and validation perplexity, as can be seen in Table 1. Moreover, the projection error seems to be constant for Dion, while for Trion the projection error has a decreasing trend for some layers, which once again shows the dynamic behavior of our column selection approach coupled with Newton-Schulz iteration.
Appendix G Pre-Training with FRUGAL/FIRA
@@
Pre-training with FRUGAL. We integrate the DCT matrix into the FRUGAL optimizer and compare against the SVD, RandPerm and Random projections. RandPerm uses a random permutation as a projection matrix, while Random generates a random, semi-orthogonal matrix. In Figure 4(a) we show the training loss for all these runs. In the zoomed-in plot for the last 5000 training steps we see that SVD projection recovers the performance full-rank AdamW, while the DCT projection is a good approximation of the SVD, which supports our claim. We present our numerical results in Table 6. We would like to emphasize the running time reduced by 1h 48m ($\approx 22.6\%$) compared to the SVD projection, as well as the memory usage reduced by about 2.2GB ($\approx 3.5\%$), while the increase in perplexity is less than one point. In comparison to RandPerm and Random projections, the runtime is on par, while train and validation perplexities are lower by approximately one point in the favor of DCT, illustrating again the benefit of our approach.
@@
Pre-training with FIRA. We integrate DCT into FIRA optimizer and compare against SVD. In Figure 4(b) we show the training loss, where our focus is on the comparison between SVD and DCT projections. We observe that DCT consistently yields lower training loss compared to the SVD projection, which is also visible in the numerical results in Table 6, translated to lower perplexities in the favor of DCT. Moreover, the memory usage is smaller by 2GB ($\approx 3\%$) and the running time is lower by 1h 54m ($\approx 23.8\%$).
@@
Appendix H Fine-Tuning
We fine-tune Llama-2 7B on GSM-8k dataset using SVD and DCT projections for FRUGAL and FIRA optimizers, as well as LD-AdamW and DCT-AdamW. We show our results for ranks $r\in\{32,512\}$ in Table 7. We also finetune Qwen-2.5-7B on GSM-8k using DCT-AdamW, GaLore and AdamW for reference and show the results in Table 8 for the same ranks.
Fine-tuning with FRUGAL. Both projections yield similar training loss for both rank values, which is an indication that DCT is indeed a good approximation for SVD. Despite having different training loss, the SVD projection achieves the same accuracy for both ranks. It is surprising that DCT yields better accuracy for lower rank compared to higher rank. In terms of memory usage, as expected for this large model, the DCT projection saves 8GB of memory ($\approx 23.4\%$) for $r=32$, while for $r=512$ the reduction in memory is only 6.3GB ($\approx 18.2\%$). The running time is reduced by roughly 35m ($\approx 75\%$) for both ranks.
Fine-tuning with FIRA. The DCT projection yields larger training loss compared to SVD. DCT recovers the accuracy, outperforms the SVD for large rank, and on average reduces the memory usage by $\approx 1.35$ GB and the running by $\approx$10m.
Fine-tuning with LDAdamW & DCT-AdamW. In this setting we compare the DCT projection with the block power-iteration used in LDAdamW as an approximation to SVD, both without EF. The training loss achieved is smaller for DCT on $r=32$ and larger for $r=512$ compared to LDAdamW. LDAdamW achieves more than 1% accuracy for $r=32$ and comparable accuracy for $r=512$. Regarding memory, DCT-AdamW uses $\approx$1 GB less memory, while achieving a speedup of about 20m for $r=512$. We would like to emphasize that EF does not help DCT-AdamW in comparison to LDAdamW for $r=32$, the accuracy of DCT-AdamW with EF being 29.11% vs 32.53% for LDAdam with EF for $r=32$, while for $r=512$ the accuracy can be recovered, scoring 35.33% compared to 35.86% for LDAdamW with EF.